What is the minimal sample size needed for a 95% confidence interval to have a maximal margin of error of 0.1 in the following scenarios? (For each answer, enter a number. Round your answers up the nearest whole number.)

(a)

a preliminary estimate for p is 0.38

(b)

there is no preliminary estimate for p
Answer #1

We need to find the minimum sample size, n, such that we have a 95% confidence interval and have a maximal margin of error of 0.1.

=> Significance level, α = 1 - 95% = 1 - 0.95 = 0.05

and Margin of error, E = 0.1

(a)

Now, when a preliminary estimate of p = 0.38 is given, the formula for sample size (n) is given by:

(b)

Now, when there is no preliminary estimate for p, the formula for sample size (n) is given by:

For any queries, feel free to comment and ask.

If the solution was helpful to you, don't forget to upvote it by clicking on the 'thumbs up' button.